OpenStudy (anonymous):

What is the domain for x in the question the number which is 3 more than x.

OpenStudy (anonymous):

We can write this out as a function. f(x)=x+3 The domain is all valid numbers for x in your function. In this case, nothing will make it invalid, or undefined since there are no fractions and x is not under a radical. So your domain is, \[-\infty \le x \le \infty \] or \[(- \infty , \infty) \]

OpenStudy (anonymous):

Wait...is the function, "3 more than x" or "3 is greater than x"

OpenStudy (anonymous):

I'm not sure either, but I interpreted in my head as you have a number y, for example, which is 3 more than x, so y=x+3.

OpenStudy (anonymous):

3 more than x

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

Ok, yeah. So you could say y=x+3. y is 3 more than x, meaning y is the larger value and x is the smaller, so you would add 3 to x as opposed to y.

OpenStudy (anonymous):

So with the equation y=x+3 or f(x)=x+3, you can find the domain by seeing which numbers would make an undefined situation. In this case there are none so its as I've shown in my first post here.

OpenStudy (anonymous):

domain is = to all of the legal inputs for x. in this equation, jlangley95 correctly identified it in his 1st post. x can be any real number.